# A Necessary Step toward Faithfulness: Measuring and Improving Consistency in Free-Text Explanations

### v1

@@ 7 Conclusion
We present a new consistency measure for free-text explanations as an important aspect of faithfulness. We optimize explanation consistency using direct preference optimization, and show that it can improve the consistency and faithfulness of model generated explanations. We hope this research advances explanation faithfulness and enhances transparency in AI-assisted decision-making.
@@ Appendix A Appendices
Model Mistral Llama-2 Yi-1.5
$k$=10 $k$=20 Avg $k$=10 $k$=20 Avg $k$=10 $k$=20 Avg
Model Mistral Llama-2 Yi-1.5
$k$=10 $k$=20 Avg $k$=10 $k$=20 Avg $k$=10 $k$=20 Avg
@@ A.1 Explanation Prompting
We use the following GPT-4 generated explanation as one-shot example to prompt models for explanations (§ 3.1) for the TripAdvisor hotel review dataset:
@@
We use the following GPT-4 generated explanation as one-shot example to prompt models for explanations (§ 3.1) for the Amazon product review dataset:
@@
A.2 Supervised Fine-tuning and Student Model Training
@@
A.3 Direct Preference Optimization
@@
We implement DPO training using TRL 0.14.0.
A.4 Explanation Faithfulness Evaluation
@@
A.5 Additional Qualitative Examples

### v2

@@ 7 Conclusion
We present a new consistency measure for free-text explanations as an important aspect of faithfulness. We optimize explanation consistency using direct preference optimization, and show that it can improve the consistency and faithfulness of model generated explanations. We hope this research advances explanation faithfulness and enhances transparency in AI-assisted decision-making. Another direction for future research is extending our approach to chain-of-thought explanations. Since our consistency measure and evaluation framework treat explanations and answers as independent variables, they are applicable regardless of how explanations are generated.
@@ A.1 Explanation Prompting
We use the following GPT-4 generated explanation as a one-shot prompt to guide models in explaining the TripAdvisor hotel review dataset (see § 3.1):
@@
To guide models in generating explanations for the Amazon product review dataset (see § 3.1), we use the following GPT-4 generated explanation as a one-shot example:
@@
A.2 Sensitivity to Prompt Format
To assess whether the PEX consistency score is sensitive to different prompt formats, we use 3 different prompt formats to compute the PEX consistency score described in Eq 4, where $\bm{e}$ is a free-text explanation to be measured:
V1: $Q^{\prime}_{1}(\bm{q},\bm{e})$ = Is this review truthful or deceptive? Review: {review}. Answer: $\bm{e}$. Therefore, this review is
V2: $Q^{\prime}_{2}(\bm{q},\bm{e})$ = Please classify the following review as either truthful or deceptive: {review}. Answer: $\bm{e}$. Hence, this review is
V3: $Q^{\prime}_{3}(\bm{q},\bm{e})$ = Evaluate the authenticity of this review: {review}. Question: Is the review truthful or deceptive? Analysis: $\bm{e}$. Answer: This review is
We then compute Kendall rank correlations between PEX consistency scores obtained from each pair of prompt versions. The results are shown in Table 1, indicating PEX scores are relatively robust to variations in prompt format.
A.3 Supervised Fine-tuning and Student Model Training
@@
Model Mistral Llama-2 Yi-1.5
kk=10 kk=20 Avg kk=10 kk=20 Avg kk=10 kk=20 Avg
Model Mistral Llama-2 Yi-1.5
kk=10 kk=20 Avg kk=10 kk=20 Avg kk=10 kk=20 Avg
A.4 Direct Preference Optimization
@@
We implement DPO training using TRL (Transformer Reinforcement Learning) 0.14.0.
A.5 Explanation Faithfulness Evaluation
@@
A.6 Additional Qualitative Examples